# Optimal High-Probability Regret for Online Convex Optimization with Two-Point Bandit Feedback

### 26 Mar 2026 v1

@@ 4 Conclusion
In this paper, we have addressed a fundamental gap in the theory of bandit online convex optimization by providing the first tight high-probability regret bound for strongly convex functions with two-point feedback. By resolving the open problem posed by Agarwal et al. (2010), we have shown that it is possible to achieve a regret bound of $O(d(\log T+\log(1/\delta))/\mu)$, which is optimal in terms of all key parameters.
Our technical contribution lies in a departure from the standard reduction-based analysis. By developing a direct analytical framework and generalizing the multi-point feedback probabilistic techniques to the high-confidence regime, we have successfully reduced the dimensional dependency from $O(d^{2})$ to $O(d)$. This result bridges the gap between expectation-based bounds and high-probability guarantees in online convex optimization settings.
@@ Appendix A Useful Lemmas
Lemma 10 (Corollary 2.6 of Ledoux (2001))
@@
Lemma 11 (Freedman’s inequality)
Let $\{Y_{t}:t\geq 0\}$ be a real-valued martingale adapted to a filtration $\{\mathcal{F}_{t}\}$, with $Y_{0}=0$. Let $Z_{t}=Y_{t}-Y_{t-1}$ denote the martingale difference sequence. Assume the sequence is uniformly bounded from above:
$Z_{t}\leq R\quad\text{a.s. for all }t\geq 1.$
Define the predictable quadratic variation process $W_{k}$ as:
$W_{t}=\sum_{j=1}^{t}\mathbb{E}[Z_{j}^{2}\mid\mathcal{F}_{j-1}].$
Then, for any $\tau\geq 0$ and $\sigma^{2}>0$:
$\mathbb{P}\left(\exists t:Y_{t}\geq\tau\text{ and }W_{t}\leq\sigma^{2}\right)\leq\exp\left(-\frac{\tau^{2}/2}{\sigma^{2}+R\tau/3}\right).$
@@
Lemma 13 (Proposition 2.6.1 of Vershynin (2026))
@@
Lemma 14 (Proposition 2.6.6 of Vershynin (2026))
@@
Lemma 15 (Proposition 2.8.1 of Vershynin (2026))
@@
Lemma 16 (Lemma 2.8.5 of Vershynin (2026))
@@
Proof First, by Lemma 10 and Eq. (15), we can obtain that $X$ is a subgaussian random variable. Furthermore, by Eq. (13), item (i) and item (iii) in Lemma 14, we can obtain that
@@
$\|\bm{g}_{t}\|^{2}\leq\frac{d^{2}}{\alpha^{2}}\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}.$ (18)
Proof Using the definition of $\bm{g}_{t}$, we have
@@
$\leq\frac{d^{2}}{2\alpha^{2}}\left[\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}\right$.
$\quad+\left.\left(\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})\right)^{2}\right]$
$=\frac{d^{2}}{\alpha^{2}}\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2},$
where the last equality is because of the symmetry of the spherical distribution.
Lemma 19
Define random variable $X_{t}=\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}\left[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})\right]$. Given $0<\delta<1$, then with a probability at least $1-\delta$, then it holds that
@@
Proof First, we introduce two random variables
$Y_{t}=X_{t}^{2},\qquad\mbox{and}\qquad V_{t}=Y_{t}-\mathbb{E}[Y_{t}].$
Define the natural filtration as the $\sigma$-algebra generated by the random variables $\bm{u}_{t}$’s observed to step $t$:
$\mathcal{F}_{t}:=\sigma(\bm{x}_{1},\bm{u}_{1},\dots,\bm{u}_{t})\mbox{ with }\mathcal{F}_{0}=\sigma(\bm{x}_{1}).$

### 05 Apr 2026 v3

@@ 4 Conclusion
In this paper, we have addressed a fundamental gap in the theory of bandit online convex optimization by providing the first tight high-probability regret bound for strongly convex functions with two-point feedback. By resolving the open problem posed by Agarwal et al., (2010), we have shown that it is possible to achieve a regret bound of $O(d(\log T+\log(1/\delta))/\mu)$, which is optimal in terms of both time horizon and dimension $d$.
Our technical contribution lies in a departure from the standard reduction-based analysis. We develop a direct analytical framework and generalizing the multi-point feedback probabilistic techniques to the high-confidence regime. Furthermore, we have successfully reduced the dimensional dependency of horizon-dependent term from $O(d^{2})$ to $O(d)$. This result bridges the gap between expectation-based bounds and high-probability guarantees in online convex optimization settings.
@@ Appendix A Useful Lemmas
Lemma 9 (Corollary 2.6 of Ledoux, (2001)).
@@
Lemma 10 (Freedman-style exponential supermartingale).
Let $(Z_{t},\mathcal{F}_{t})$ be a martingale difference sequence such that $Z_{t}\leq b$ almost surely for all $t$, where $b>0$ is a constant. Define
$V_{t}:=\sum_{s=1}^{t}\mathbb{E}[Z_{s}^{2}\mid\mathcal{F}_{s-1}].$
Then for every $\lambda\in[0,3/b)$, the process
$M_{t}(\lambda):=\exp\!\left(\lambda\sum_{s=1}^{t}Z_{s}-\frac{\lambda^{2}}{2(1-\lambda b/3)}V_{t}\right)$
is a nonnegative supermartingale. Consequently, for every $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sum_{t=1}^{T}Z_{t}\leq\frac{\lambda}{2(1-\lambda b/3)}V_{T}+\frac{1}{\lambda}\log\frac{1}{\delta}.$
Proof.
Fix $\lambda\geq 0$. For every $x\leq b$,
$e^{\lambda x}=1+\lambda x+\sum_{k=2}^{\infty}\frac{\lambda^{k}x^{k}}{k!}.$
Since $x\leq b$, we have $x^{k}\leq x^{2}b^{k-2}$ for all $k\geq 2$. Hence
$e^{\lambda x}\leq 1+\lambda x+\frac{x^{2}}{b^{2}}\sum_{k=2}^{\infty}\frac{(\lambda b)^{k}}{k!}=1+\lambda x+\frac{\phi(\lambda b)}{b^{2}}x^{2}.$
Applying this with $x=Z_{t}$ and taking conditional expectation given $\mathcal{F}_{t-1}$,
$\mathbb{E}[e^{\lambda Z_{t}}\mid\mathcal{F}_{t-1}]\leq 1+\frac{\phi(\lambda b)}{b^{2}}\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}],$
where we used $\mathbb{E}[Z_{t}\mid\mathcal{F}_{t-1}]=0$. Using $1+u\leq e^{u}$ gives
$\mathbb{E}[e^{\lambda Z_{t}}\mid\mathcal{F}_{t-1}]\leq\exp\!\left(\frac{\phi(\lambda b)}{b^{2}}\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}]\right).$
Thus,
$\mathbb{E}\!\left[\exp\!\left(\lambda Z_{t}-\frac{\phi(\lambda b)}{b^{2}}\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}]\right)\Bigm|\mathcal{F}_{t-1}\right]\leq 1.$
The random process
$\widetilde{M}_{t}(\lambda):=\exp\!\left(\lambda\sum_{s=1}^{t}Z_{s}-\frac{\phi(\lambda b)}{b^{2}}V_{t}\right),$
which can be represented as
$\widetilde{M}_{t}(\lambda)=\widetilde{M}_{t-1}(\lambda)\exp\!\left(\lambda Z_{t}-\frac{\phi(\lambda b)}{b^{2}}\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}]\right),$
satisfies
$\mathbb{E}[\widetilde{M}_{t}(\lambda)\mid\mathcal{F}_{t-1}]\leq\widetilde{M}_{t-1}(\lambda).$
So $(\widetilde{M}_{t}(\lambda))_{t\geq 0}$ is a nonnegative supermartingale.
Finally, for $0\leq u<3$,
$\phi(u)=\sum_{k=2}^{\infty}\frac{u^{k}}{k!}\leq\sum_{k=2}^{\infty}\frac{u^{k}}{2\cdot 3^{k-2}}=\frac{u^{2}}{2(1-u/3)},$
because $k!\geq 2\cdot 3^{k-2}$ for all $k\geq 2$. Substituting $u=\lambda b$ shows that for $\lambda\in[0,3/b)$,
$\frac{\phi(\lambda b)}{b^{2}}\leq\frac{\lambda^{2}}{2(1-\lambda b/3)}.$
Hence
$M_{t}(\lambda)=\exp\!\left(\lambda\sum_{s=1}^{t}Z_{s}-\frac{\lambda^{2}}{2(1-\lambda b/3)}V_{t}\right)\leq\widetilde{M}_{t}(\lambda),$
and therefore $(M_{t}(\lambda))_{t\geq 0}$ is also a nonnegative supermartingale.
Applying Markov’s inequality, for any $\delta\in(0,1)$,
$\Pr\!\left(\lambda\sum_{t=1}^{T}Z_{t}-\frac{\lambda^{2}}{2(1-\lambda b/3)}V_{t}\geq\log\frac{1}{\delta}\right)$ $=\Pr\!\left(M_{T}(\lambda)\geq\frac{1}{\delta}\right)$
$\leq\delta\,\mathbb{E}[M_{T}(\lambda)]\leq\delta.$
Equivalently, with probability at least $1-\delta$,
$\sum_{t=1}^{T}Z_{t}\leq\frac{\lambda}{2(1-\lambda b/3)}V_{T}+\frac{1}{\lambda}\log\frac{1}{\delta}.$
@@
Lemma 11 (Proposition 2.6.1 of Vershynin, (2026)).
@@
Lemma 12 (Proposition 2.6.6 of Vershynin, (2026)).
@@
Lemma 13 (Proposition 2.8.1 of Vershynin, (2026)).
@@
Lemma 14 (Lemma 2.8.5 of Vershynin, (2026)).
@@
Proof.
First, by Lemma 9 and Eq. (15), we can obtain that $X$ is a subgaussian random variable. Furthermore, by Eq. (13), item (i) and item (iii) in Lemma 12, we can obtain that
@@
$\|\bm{g}_{t}\|^{2}\leq\frac{3d^{2}}{4\alpha^{2}}\left(\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}+\left(\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})]-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})\right)^{2}\right).$ (18)
Proof.
Using the definition of $\bm{g}_{t}$, we have
@@
$\leq\frac{3d^{2}}{4\alpha^{2}}\left[\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}\right$.
$\quad+\left(\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})]\right)^{2}+\left.\left(\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})]-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})\right)^{2}\right]$
$=\frac{3d^{2}}{4\alpha^{2}}\left(\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}+\left(\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})]-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})\right)^{2}\right),$
where the last equality is because of the symmetry of the spherical distribution. ∎
Lemma 17.
Define the natural filtration $\mathcal{F}_{t}:=\sigma(\bm{x}_{1},\bm{u}_{1},\dots,\bm{u}_{t})$ with $\mathcal{F}_{0}=\sigma(\bm{x}_{1})$ as the $\sigma$-algebra generated by the random variables $\bm{u}_{t}$’s observed to step $t$. Define random variable $X_{t}=\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}\left[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})\mid\mathcal{F}_{t-1}\right]$. Given $0<\delta<1$, then with a probability at least $1-\delta$, then it holds that
@@
Proof.
First, we introduce two random variables
$Y_{t}=X_{t}^{2},\qquad\mbox{and}\qquad V_{t}=Y_{t}-\mathbb{E}[Y_{t}\mid\mathcal{F}_{t}].$